OpenStudy (anonymous):

You roll two die. What is the probability that that the sum of the numbers with be a multiple of 5. EXPLAIN.

OpenStudy (anonymous):

This is probably more suited for the mathematics category :P BUT. Here are the possible cominations: 1,1 =2 1,2 =3 1,3 =4 1,4 =5 - 1,5 =6 1,6 =7 2,1 =3 2,2 =4 2,3 =5 - 2,4 =6 2,5 =7 2,6 =8 3,1 =4 3,2 =5 - 3,3 =6 3,4 =7 3,5 =8 3,6 =9 4,1 =5 - 4,2 =6 4,3 =7 4,4 =8 4,5 =9 4,6 =10 - 5,1 =6 5,2 =7 5,3 =8 5,4 =9 5,5 =10 - 5,6 =11 6,1 =7 6,2 =8 6,3 =9 6,4 =10 - 6,5 =11 6,6 =12 There are 36 possible combinations and 7 possible combinations resulting in a sum of a number that is a multiple of 5. Thus the probability is 7/36

OpenStudy (anonymous):

i know it was an accident

OpenStudy (anonymous):

Hahaha its all good :)